Circular No. 6197

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995U IN ESO 235-IG13 B. P. Schmidt, Mount Stromlo and Siding Spring Observatories, re-

ports the discovery of a supernova of mag ∼ 18 by R. H. McNaught on a U.K. Schmidt R film taken by M. J. Drinkwater on July 29.56 UT. SN 1995U is located at α = 20^{h}54^{m}40^{s}.62, δ = −51^{o}22′26′′.8 (equinox 2000.0) in the galaxy ESO 235-IG13. A nearby star of mag 17–17.5 has end figures 41^{s}.19, 15′′.1. M. M. Phillips, Cerro Tololo Interamerican Observatory, reports that BV I CCD images obtained by P. Harding with the CTIO 0.9-m telescope on Aug. 1 and 2 clearly confirm the presence of the supernova and indicate that it decreased in brightness by ∼ 0.1 mag in both B and V between the two nights.

SUPERNOVA 1995V IN NGC 1087 R. O. Evans, Coonabarabran, N.S.W., reports his visual discovery of a

supernova in NGC 1087 (α = 2^{h}46^{m}.4, δ = −0^{o}30′, equinox 2000.0) with the Australian National University’s 1-m telescope at Siding Spring (J. Jarman and T. Cragg assisting) on Aug. 1.8 UT. An unfiltered CCD image obtained by M. Dopita and C. Trung Hua with the 2.3-m reflector confirms SN 1995V at mag 15.0 and 21′′ east and 3′′ south of the galaxy’s center. It is likely that the object was not present when Evans examined the galaxy on July 25 with a 0.41-m reflector. S. Benetti, European Southern Observatory, reports that inspection of a preliminarily-reduced CCD spectrum (range 396–793 nm, resolution 1 nm) obtained on Aug. 2.40 UT with the Danish 1.54-m telescope + DFOSC at La Silla confirms the object as a type II supernova around maximum. Strong Hγ, Hβ and Hα lines with P-Cyg profiles are superimposed on a blue continuum; the expansion velocity deduced from their minima is ∼ 11 000 km/s (the recession velocity of NGC 1087 being 1519 km/s). The expansion velocity deduced from the FWHM of Hα is ∼ 15 500 km/s. The He I 587.6-nm line is also present with a P-Cyg profile, and it shows a somewhat slower expansion velocity (∼ 9000 km/s).

COMET 6P/D’ARREST Visual magnitudes: July 25.40 UT, 9.7 (A. Hale, Cloudcroft, NM, 0.41-

m reflector); 26.35, 8.8 (R. Keen, Mt. Thorodin, CO, 7 × 50 binoculars); 26.95, 9.4 (M. Möller, Timmendorfer Strand, Germany, 0.15-m reflector); 29.38, 9.6 (Hale); 30.33, 8.7 (C. S. Morris, Pine Mountain Club, CA, 20 × 80 binoculars); 31.04, 8.5 (K. Hornoch, Lelekovice, Czech Rep., 0.20-m reflector); 31.65, 8.4 (T. Lovejoy, Brisbane, Australia, 15 × 80 binoculars).

1995 August 2 Brian G. Marsden